# Learning Distributions on Manifolds with Free-form Flows

### v1

@@ 6 Conclusion
In this paper, we present manifold free-form flows (M-FFF) to learn a probability distribution on a manifold. They are based on a neural network construction that learns a function in an embedding space whose outputs are projected to the manifold (Fig. 1). This naturally makes the learned distributions adhere to the topology of the manifold at hand.
The distribution on the manifold is then learned by adapting the framework of free-form flows (FFF) to manifolds. We extend it to estimate the gradient of the change of variables formula for manifold-to-manifold functions.
The main advantage of M-FFF is that they offer competitive performance similar to previous ODE- or SDE-based methods at greatly reduced sampling costs, typically between two and three orders of magnitude faster. In addition, they are very easy to set up, since only a projection from the embedding space to the manifold needs to be known.
@@ Acknowledgements
Supplementary Material
7 Free-form flows on Riemannian manifolds
In this appendix, we will focus on intuitive definitions of concepts from topology and differential geometry. For a more rigorous treatment of these concepts, see [22].
@@ 7 Free-form flows on Riemannian manifolds
7.1 Manifold change of variables
@@
$\operatorname{proj}:P\rightarrow\mathbb{R}^{m}$ (21)
where $P\subseteq\mathbb{R}^{m}$ is the projectable set. We require the projection to have the following properties (the first is true of all projections, the others are additional requirements):
@@
    $\operatorname{proj}\circ\operatorname{proj}=\operatorname{proj}$
@@
    $\operatorname{rank}(\operatorname{proj}^{\prime}(\operatorname{proj}(x)))=n$ for all $x\in P$
@@
=$\mathcal{M}=\{x\in\mathbb{R}^{m}:\operatorname{proj}(x)=x\}$ (22)
@@
$\mathcal{T}_{x}\mathcal{M}=\operatorname{col}(\operatorname{proj}^{\prime}(x))$ (23)
where $\operatorname{col}$ denotes the column space. Since the rank of $\operatorname{proj}^{\prime}(x)$ with $x\in\mathcal{M}$ is $n$, the tangent space is $n$-dimensional and $\mathcal{M}$ is an $n$-dimensional manifold. The Riemannian metric $g_{x}:\mathcal{T}_{x}\mathcal{M}\times\mathcal{T}_{x}\mathcal{M}\rightarrow\mathbb{R}$ defined on this manifold is realized as an $m\times m$ matrix $G(x)$. If $\mathcal{M}$ is isometrically embedded then $G(x)$ is just the identity matrix.
The Jacobian of the projection is a projection matrix, meaning $\operatorname{proj}^{\prime}(x)\operatorname{proj}^{\prime}(x)=\operatorname{proj}^{\prime}(x)$ for $x\in\mathcal{M}$. For any $v$ in the column space of $\operatorname{proj}^{\prime}(x)$, there is a $u$ such that $v=\operatorname{proj}^{\prime}(x)u$ and due to the projection property, $\operatorname{proj}^{\prime}(x)v=\operatorname{proj}^{\prime}(x)u=v$. Similarly, for any $w$ in the row space of $\operatorname{proj}^{\prime}(x)$, $w\operatorname{proj}^{\prime}(x)=w$. If $\operatorname{proj}$ is an orthogonal projection, $\operatorname{proj}^{\prime}$ is symmetric by definition and hence the row and column spaces are identical.
@@
where $G(\phi(u))$ is the matrix representation of $g_{\phi(u))}$. The integral on the right is an ordinary integral in $\mathbb{R}^{n}$. The quantity inside the determinant is known as the pullback metric.
@@
Let $(\mathcal{M}_{X},g_{X})$ and $(\mathcal{M}_{Z},g_{Z})$ be $n$-dimensional Riemannian manifolds embedded in $\mathbb{R}^{m}$ with the same global topological structure. Let $p_{X}$ be a probability distribution on $\mathcal{M}_{X}$ and let $f:\mathcal{M}_{X}\rightarrow\mathcal{M}_{Z}$ be a diffeomorphism. Let $p_{Z}$ be the pushforward of $p_{X}$ under $f$ (i.e. if $p_{X}$ is the probability density of $X$, $p_{Z}$ is the probability density of $f(X)$).
Let $x\in\mathcal{M}_{X}$. Let $Q\in\mathbb{R}^{m\times n}$ be an orthonormal basis for $\mathcal{T}_{x}\mathcal{M}_{X}$ and $R\in\mathbb{R}^{m\times n}$ an orthonormal basis for $\mathcal{T}_{f(x)}\mathcal{M}_{Z}$.
Then $p_{X}$ and $p_{Z}$ are related under the change of variables $x\mapsto f(x)$ by:
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|R^{T}f^{\prime}(x)Q|\\ +\tfrac{1}{2}\log\frac{|R^{T}G_{Z}(f(x))R|}{|Q^{T}G_{X}(x)Q|}.$ (25)
Note that $Q$ and $R$ depend on $x$ and $f(x)$ respectively, but this is omitted for brevity.
In the main text we state a simplified form of the equation:
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|f^{\prime}(x)|+\tfrac{1}{2}\log\frac{|G_{Z}(f(x))|}{|G_{X}(x)|}$ (26)
with the statement that the determinants have to be evaluated in the tangent spaces, not in the ambient $m$-dimensional space. This is equivalent to the statement of Theorem 1.
@@
Let $\phi:\mathbb{R}^{n}\rightarrow\mathcal{M}_{X}$ be defined by $\phi(u)=\operatorname{proj}_{X}(x+Qu)$. Let $U$ be an open subset of $\mathbb{R}^{n}$ containing the origin which is small enough so that $\phi$ is bijective. Let $\psi:\mathbb{R}^{n}\rightarrow\mathcal{M}_{Z}$ be defined by $\psi(w)=\operatorname{proj}_{Z}(f(x)+Rw)$. Define $\varphi=\psi^{-1}\circ f\circ\phi$ and let $W=\varphi(U)$.
Note that $\phi^{\prime}(u)=\operatorname{proj}_{X}^{\prime}(x+Qu)\cdot Q$ and hence $\phi^{\prime}(0)=\operatorname{proj}_{X}^{\prime}(x)Q=Q$ (since each column of $Q$ is in $\mathcal{T}_{x}\mathcal{M}_{X}=\operatorname{col}(\operatorname{proj}_{X}^{\prime}(x))$).
Similarly, $\psi^{\prime}(0)=R$. Since $\psi$ is a map from $n$ to $m$ dimensions, there is not a unique function from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$ which is $\psi^{-1}$ on the manifold and there are remaining degrees of freedom which can result in different Jacobians. For our purposes, we define the inverse $\psi^{-1}$ such that $\psi\circ\psi^{-1}$ is an orthogonal projection onto $\mathcal{M}_{Z}$. This means $\psi^{\prime}(\psi^{-1}(f(x)))(\psi^{-1})^{\prime}(f(x))=RR^{T}$ and hence $(\psi^{-1})^{\prime}(f(x))=R^{T}$.
@@
Note that $Q$ and $R$ also depend on $u$. Changing variables of the RHS with $w=\varphi(u)$ gives us
@@
$p_{X}(x)\sqrt{|Q^{T}G_{X}(x)Q|}\\ =p_{Z}(f(x))\sqrt{|R^{T}G_{Z}(f(x))R|}\cdot\left|\frac{\partial w}{\partial u}\right|.$ (30)
where now $Q$ is evaluated at $u=0$ and $R$ at $w=\varphi(0)$. Since $w=\psi^{-1}(f(\phi(u)))$, the Jacobian has the following form when evaluated at the origin (note $\phi(0)=x$):
@@
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|R^{T}f^{\prime}(x)Q|\\ +\tfrac{1}{2}\log\frac{|R^{T}G_{Z}(f(x))R|}{|Q^{T}G_{X}(x)Q|}.$ (33)
@@
Since the probability mass in both regions should be the same we can write
$p_{X}(x)\operatorname{vol}(U)\approx p_{Z}(z)\operatorname{vol}(f(U))$ (39)
@@
7.2 Loss function
In this section, we consider $f_{\theta}$ parameterized by $\theta$, but drop the subscript for ease of presentation.
@@ 7.2 Loss function
Let the assumptions of Theorem 1 hold and let $g:\mathcal{M}_{Z}\rightarrow\mathcal{M}_{X}$ be the inverse of $f$. Then, the derivative of the change of variables term has the following trace expression:
$\nabla_{\theta}\log|R^{T}f^{\prime}(x)Q|=\operatorname{tr}(R^{T}(\nabla_{\theta}f^{\prime}(x))g^{\prime}(f(x))R).$ (41)
@@
First, a reminder that $\varphi^{\prime}(u)=R^{T}f^{\prime}(x)Q$ with $\varphi=\psi^{-1}\circ f\circ\phi$. Let $\chi=\varphi^{-1}$, i.e. $\chi=\phi^{-1}\circ g\circ\psi$. Jacobi’s formula tells us that
@@
$\nabla_{\theta}\log|R^{T}f^{\prime}(x)Q|=\operatorname{tr}(\nabla_{\theta}(R^{T})f^{\prime}(x)QQ^{T}g^{\prime}(f(x))R)\\ +\operatorname{tr}(R^{T}\nabla_{\theta}(f^{\prime}(x))QQ^{T}g^{\prime}(f(x))R).$ (47)
Since $Q$ is an orthonormal basis for $\mathcal{T}_{x}\mathcal{M}_{X}$, $QQ^{T}$ is a projection matrix onto $\mathcal{T}_{x}\mathcal{M}_{X}$. This is because $(QQ^{T})^{2}=QQ^{T}QQ^{T}=QQ^{T}$, using $Q^{T}Q=I$. As a result, $QQ^{T}\operatorname{proj}^{\prime}(x)=\operatorname{proj}^{\prime}(x)$. Since $g$ is wrapped inside a projection: $g(z)=\operatorname{proj}_{Z}(\tilde{g}(z))$, therefore $g^{\prime}(z)=\operatorname{proj}_{Z}^{\prime}(\tilde{g}(z))\tilde{g}^{\prime}(z)$, so $QQ^{T}g^{\prime}(z)=g^{\prime}(z)$. Note also that $f^{\prime}(x)g^{\prime}(f(x))=I$ since $g$ is the inverse of $f$. This simplifies the equation:
$\nabla_{\theta}\log|R^{T}f^{\prime}(x)Q|=\operatorname{tr}(\nabla_{\theta}(R^{T})R)\\ +\operatorname{tr}(R^{T}\nabla_{\theta}(f^{\prime}(x))g^{\prime}(f(x))R)$ (48)
@@
In the above proof we used the fact that $QQ^{T}g^{\prime}(z)=g^{\prime}(z)$. Can we use $RR^{T}f^{\prime}(x)=f^{\prime}(x)$ to simplify the equation further? No, we cannot, since the expression involving $f^{\prime}$ is actually its derivative with respect to parameters, which may not have the same matrix structure as $f^{\prime}$. Is it instead possible to use $g^{\prime}(z)RR^{T}=g^{\prime}(z)$ for simplification? If $g$ is implemented as $\operatorname{proj}_{Z}(\tilde{g}(z))$, this is not necessarily true, as $g^{\prime}(z)$ might not be a map from the tangent space at $z$ to the tangent space at $g(z)$. For example, if we add a small deviation $v$ to $z$, where $v$ is orthogonal to the tangent space at $z$, then $g(z+v)$ might not equal $g(z)$. However, this would mean that derivatives in the off-manifold direction can be non-zero, meaning that $g^{\prime}(z)v\neq g^{\prime}(z)RR^{T}v=0$ (since $RR^{T}$ will project $v$ to 0). We can change this by prepending $g$ by a projection:
$g=\operatorname{proj}_{X}\circ\tilde{g}\circ\operatorname{proj}_{Z}.$ (50)
If $\operatorname{proj}_{Z}$ is an orthogonal projection, meaning that $\operatorname{proj}_{Z}^{\prime}$ is symmetric, the column space and row space of $\operatorname{proj}_{Z}$ will both be the same as those of $RR^{T}$, meaning $\operatorname{proj}_{Z}^{\prime}(z)RR^{T}=\operatorname{proj}_{Z}^{\prime}$ and hence $g^{\prime}(z)RR^{T}=g^{\prime}(z)$. This leads to the following corollary:
@@
Suppose the assumptions of Theorem 1 hold, and $g$ is the inverse of $f$ with the following implementation:
$g=\operatorname{proj}_{X}\circ\tilde{g}\circ\operatorname{proj}_{Z}$ (51)
where $\operatorname{proj}_{Z}$ is an orthogonal projection. Then the derivative of the change of variables term has the following trace expression:
$\nabla_{\theta}\log|R^{T}f^{\prime}(x)Q|=\operatorname{tr}((\nabla_{\theta}f^{\prime}(x))g^{\prime}(f(x))).$ (52)
@@
Take the result of Theorem 2 and use the cyclic property of the trace and the properties of $g^{\prime}$ discussed above:
@@
We have two variants of the trace estimate derived above, one evaluated in $\mathbb{R}^{n}$, the other in $\mathbb{R}^{m}$. The first can be estimated using the following equality:
@@
$\operatorname{tr}(\nabla_{\theta}(f^{\prime}(x))g^{\prime}(f(x)))\\ =\nabla_{\theta}\mathbb{E}_{p(v)}[v^{T}\nabla_{\theta}(f^{\prime}(x))\texttt{SG}[g^{\prime}(f(x))]v].$ (61)
7.3 Variance reduction
When using a Hutchinson trace estimator with standard normal $v\in R^{n}$, we can reduce the variance of the estimate by scaling $v$ to have length $\sqrt{n}$ (see [18]). The scaled variable will still have zero mean and unit covariance so the estimate remains unbiased, but the variance is reduced, with the effect especially pronounced in low dimensions.

### v2

@@ 6 Conclusion
In this paper, we present Manifold Free-Form Flows (M-FFF), a generative model designed for manifold data. To the best of our knowledge, it is the first generative model on manifolds with single-step sampling and density estimation readily applicable to arbitrary Riemannian manifolds. This significantly accelerates inference and allows for deployment on edge devices.
M-FFF matches or outperforms single-step architectures specialized to particular manifolds. It also surpasses multi-step methods in some cases, despite the greatly reduced inference compute.
Adapting M-FFF to new manifolds is straightforward and only requires selecting an embedding space and a projection to the manifold. In contrast, competing multi-step methods are more challenging to adapt as they require implementing a diffusion process or computing distances on the manifold.
@@ Acknowledgements
Appendix A Free-form flows on Riemannian manifolds
In this appendix, we will focus on intuitive definitions of concepts from topology and differential geometry. For a more rigorous treatment of these concepts, see (Jost, 2008).
@@ Appendix A Free-form flows on Riemannian manifolds
A.1 Manifold change of variables
@@
$\pi:\mathbb{P}\rightarrow\mathbb{R}^{m}$ (23)
where $\mathbb{P}\subseteq\mathbb{R}^{m}$ is the projectable set. We require the projection to have the following properties (the first is true of all projections, the others are additional requirements):
@@
    $\pi\circ\pi=\pi$
@@
    $\operatorname{rank}(\pi^{\prime}(\pi(x)))=n$ for all $x\in\mathbb{P}$
@@
=$\mathcal{M}=\{x\in\mathbb{R}^{m}:\pi(x)=x\}$ (24)
@@
$\mathcal{T}_{x}\mathcal{M}=\operatorname{col}(\pi^{\prime}(x))$ (25)
where $\operatorname{col}$ denotes the column space. Since the rank of $\pi^{\prime}(x)$ with $x\in\mathcal{M}$ is $n$, the tangent space is $n$-dimensional and $\mathcal{M}$ is an $n$-dimensional manifold. To avoid clutter we denote the Riemannian metric and its $m\times m$ matrix representation with $G$ interchangeably. If $\mathcal{M}$ is isometrically embedded then $G(x)$ is just the identity matrix.
The Jacobian of the projection is a projection matrix, meaning $\pi^{\prime}(x)\pi^{\prime}(x)=\pi^{\prime}(x)$ for $x\in\mathcal{M}$. For any $v$ in the column space of $\pi^{\prime}(x)$, there is a $u$ such that $v=\pi^{\prime}(x)u$ and due to the projection property, $\pi^{\prime}(x)v=\pi^{\prime}(x)u=v$. Similarly, for any $w$ in the row space of $\pi^{\prime}(x)$, $w\pi^{\prime}(x)=w$. If $\pi$ is an orthogonal projection, $\pi^{\prime}$ is symmetric by definition and hence the row and column spaces are identical.
@@
The integral on the right is an ordinary integral in $\mathbb{R}^{n}$. The quantity inside the determinant is known as the pullback metric.
@@
Let $(\mathcal{M}_{X},G_{X})$ and $(\mathcal{M}_{Z},G_{Z})$ be $n$-dimensional Riemannian manifolds embedded in $\mathbb{R}^{m}$, i.e., $\mathcal{M}_{X},\mathcal{M}_{Z}\subseteq\mathbb{R}^{m}$, and assume they have the same global topological structure. Let $p_{X}$ be a probability distribution on $\mathcal{M}_{X}$ and let $f:\mathcal{M}_{X}\rightarrow\mathcal{M}_{Z}$ be a diffeomorphism. Let $p_{Z}$ be the pushforward of $p_{X}$ under $f$ (i.e., if $p_{X}$ is the probability density of $X$, then $p_{Z}$ is the probability density of $f(X)$).
Let $x\in\mathcal{M}_{X}$. Define $Q\in\mathbb{R}^{m\times n}$ as an orthonormal basis for $\mathcal{T}_{x}\mathcal{M}_{X}$ and $R\in\mathbb{R}^{m\times n}$ as an orthonormal basis for $\mathcal{T}_{f(x)}\mathcal{M}_{Z}$.
Then, the probability densities $p_{X}$ and $p_{Z}$ are related under the change of variables $x\mapsto f(x)$ by the following equation:
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|R^{T}f^{\prime}(x)Q|+\tfrac{1}{2}\log\frac{|R^{T}G_{Z}(f(x))R|}{|Q^{T}G_{X}(x)Q|}.$ (27)
where $Q$ and $R$ depend on $x$ and $f(x)$, respectively, although this dependency is omitted for brevity.
@@
Let $\phi:\mathbb{R}^{n}\rightarrow\mathcal{M}_{X}$ be defined by $\phi(u)=\pi_{X}(x+Qu)$. Let $U$ be an open subset of $\mathbb{R}^{n}$ containing the origin which is small enough so that $\phi$ is bijective. Let $\psi:\mathbb{R}^{n}\rightarrow\mathcal{M}_{Z}$ be defined by $\psi(w)=\pi_{Z}(f(x)+Rw)$. Define $\varphi=\psi^{-1}\circ f\circ\phi$ and let $W=\varphi(U)$.
Note that $\phi^{\prime}(u)=\pi_{X}^{\prime}(x+Qu)\cdot Q$ and hence $\phi^{\prime}(0)=\pi_{X}^{\prime}(x)Q=Q$ (since each column of $Q$ is in $\mathcal{T}_{x}\mathcal{M}_{X}=\operatorname{col}(\pi_{X}^{\prime}(x))$).
Similarly, $\psi^{\prime}(0)=R$. Since $\psi$ is a map from $n$ to $m$ dimensions, there is not a unique function from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$ which is $\psi^{-1}$ on the manifold and there are remaining degrees of freedom in the off-manifold behavior which can result in different Jacobians. For our purposes, we define the inverse $\psi^{-1}$ such that $\psi\circ\psi^{-1}$ is an orthogonal projection onto $\mathcal{M}_{Z}$. This means $\psi^{\prime}(\psi^{-1}(f(x)))(\psi^{-1})^{\prime}(f(x))=RR^{T}$ and hence $(\psi^{-1})^{\prime}(f(x))=R^{T}$.
@@
Changing variables of the RHS with $w=\varphi(u)$ gives us
@@
$p_{X}(x)\sqrt{|Q^{T}G_{X}(x)Q|}=p_{Z}(f(x))\sqrt{|R^{T}G_{Z}(f(x))R|}\cdot\left|\frac{\partial w}{\partial u}\right|.$ (31)
Since $w=\psi^{-1}(f(\phi(u)))$, the Jacobian has the following form when evaluated at the origin (note $\phi(0)=x$):
@@
$\log p_{X}(x)=\log p_{Z}(f(x))+\log|R^{T}f^{\prime}(x)Q|+\tfrac{1}{2}\log\frac{|R^{T}G_{Z}(f(x))R|}{|Q^{T}G_{X}(x)Q|}.$ (34)
@@
The probability density in $U$ and $f(U)$ should be roughly constant since both regions are very small. Since the probability mass in both regions should be the same we can write
$p_{X}(x)\operatorname{vol}(U)\approx p_{Z}(f(x))\operatorname{vol}(f(U))$ (40)
@@
A.2 Loss function
@@
Under the assumptions of Theorem 2 with $f=f_{\theta}$. Let $v\in\mathbb{R}^{m}$ be a random variable with zero mean and covariance $RR^{T}$. Then, the derivative of the change of variables term has the following trace expression, where $z=f_{\theta}(x)$:
$\nabla_{\theta}\log|R^{T}f_{\theta}^{\prime}(x)Q|$ $=\operatorname{tr}(R^{T}(\nabla_{\theta}f_{\theta}^{\prime}(x)){f_{\theta}^{-1}}^{\prime}(z)R)$ (42)
$=\mathbb{E}_{v}\left[v^{T}(\nabla_{\theta}f_{\theta}^{\prime}(x)){f_{\theta}^{-1}}^{\prime}(z)v\right].$ (43)
@@
For brevity, we drop the index $\theta$ and denote $g=f^{-1}$. First, a reminder that $\varphi^{\prime}(u)=R^{T}f^{\prime}(x)Q$ with $\varphi=\psi^{-1}\circ f\circ\phi$. Let $\chi=\varphi^{-1}$, i.e. $\chi=\phi^{-1}\circ g\circ\psi$. Jacobi’s formula tells us that
@@
$\nabla_{\theta}\log|R^{T}f^{\prime}(x)Q|=\operatorname{tr}(\nabla_{\theta}(R^{T})f^{\prime}(x)QQ^{T}g^{\prime}(f(x))R)+\operatorname{tr}(R^{T}\nabla_{\theta}(f^{\prime}(x))QQ^{T}g^{\prime}(f(x))R).$ (49)
Since $Q$ is an orthonormal basis for $\mathcal{T}_{x}\mathcal{M}_{X}$, $QQ^{T}$ is a projection matrix onto $\mathcal{T}_{x}\mathcal{M}_{X}$. This is because $(QQ^{T})^{2}=QQ^{T}QQ^{T}=QQ^{T}$, using $Q^{T}Q=I$. As a result, $QQ^{T}\pi^{\prime}(x)=\pi^{\prime}(x)$. Since $g$ can also be written inside a projection: $g(z)=\pi_{Z}(g(z))$, therefore $g^{\prime}(z)=\pi_{Z}^{\prime}(\tilde{g}(z))\tilde{g}^{\prime}(z)$, so $QQ^{T}g^{\prime}(z)=g^{\prime}(z)$. Note also that $f^{\prime}(x)g^{\prime}(f(x))=I$ since $f\circ g=\operatorname{id}$. This simplifies the equation:
$\nabla_{\theta}\log|R^{T}f^{\prime}(x)Q|=\operatorname{tr}(\nabla_{\theta}(R^{T})R)+\operatorname{tr}(R^{T}\nabla_{\theta}(f^{\prime}(x))g^{\prime}(f(x))R)$ (50)
@@
In the above proof we used the fact that $QQ^{T}g^{\prime}(z)=g^{\prime}(z)$, where we dropped the index $\theta$ and use $g:=f^{-1}$ for brevity. Can we use $RR^{T}f^{\prime}(x)=f^{\prime}(x)$ to simplify the equation further? No, we cannot, since the expression involving $f^{\prime}$ is actually its derivative with respect to parameters, which may not have the same matrix structure as $f^{\prime}$. Is it instead possible to use $g^{\prime}(z)RR^{T}=g^{\prime}(z)$ for simplification? If $g$ is implemented as $\pi_{Z}(\tilde{g}(z))$, this is not necessarily true, as $g^{\prime}(z)$ might not be a map from the tangent space at $z$ to the tangent space at $g(z)$. For example, if we add a small deviation $v$ to $z$, where $v$ is orthogonal to the tangent space at $z$, then $g(z+v)$ might not equal $g(z)$. However, this would mean that derivatives in the off-manifold direction can be non-zero, meaning that $g^{\prime}(z)v\neq g^{\prime}(z)RR^{T}v=0$ (since $RR^{T}$ will project $v$ to 0). We can change this by prepending $g$ by a projection:
$g=\pi_{X}\circ\tilde{g}\circ\pi_{Z}.$ (52)
If $\pi_{Z}$ is an orthogonal projection, meaning that $\pi_{Z}^{\prime}$ is symmetric, the column space and row space of $\pi_{Z}$ will both be the same as those of $RR^{T}$, meaning $\pi_{Z}^{\prime}(z)RR^{T}=\pi_{Z}^{\prime}$ and hence $g^{\prime}(z)RR^{T}=g^{\prime}(z)$. This leads to the following corollary:
@@
Suppose the assumptions of Theorem 2 hold with $f=f_{\theta}$ and the following implementation:
$f_{\theta}^{-1}=\pi_{X}\circ f_{\theta}^{-1}\circ\pi_{Z}$ (53)
where $\pi_{Z}$ is an orthogonal projection. Then the derivative of the change of variables term has the following trace expression, where $z=f_{\theta}(x)$:
$\nabla_{\theta}\log|R^{T}f_{\theta}^{\prime}(x)Q|=\operatorname{tr}((\nabla_{\theta}f^{\prime}(x))(f_{\theta}^{-1})^{\prime}(z)).$ (54)
@@
Again, we drop the index $\theta$ and let $g=f^{-1}$ for brevity. Take the result of Theorem 3 and use the cyclic property of the trace and the properties of $g^{\prime}$ discussed above:
@@
We have two variants of the trace estimate derived above, one evaluated in $\mathbb{R}^{n}$, the other in $\mathbb{R}^{m}$. The first can be estimated using the following equality (again dropping the index $\theta$ and using $g=f^{-1}$):
@@
$\operatorname{tr}(\nabla_{\theta}(f^{\prime}(x))g^{\prime}(f(x)))=\nabla_{\theta}\mathbb{E}_{p(v)}[v^{T}\nabla_{\theta}(f^{\prime}(x))\texttt{SG}[g^{\prime}(f(x))]v].$ (63)
A.3 Variance reduction
When using a Hutchinson trace estimator with standard normal $v\in R^{n}$, we can reduce the variance of the estimate by scaling $v$ to have length $\sqrt{n}$ (see (Girard, 1989)). The scaled variable will still have zero mean and unit covariance so the estimate remains unbiased, but the variance is reduced, with the effect especially pronounced in low dimensions.